OpenStudy (anonymous):

Which correctly describes how the graph of the inequality 4x - 2y > 20 is shaded?

OpenStudy (tkhunny):

Try the point (0,0). 4(0) - 2(0) = 0 < 20 -- That doesn't work! Okay, so (0,0) better not be shaded!

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

oh sorry forgot A. Above the solid line B. Below the solid line C. Above the dashed line D. Under the dashed line

OpenStudy (anonymous):

can you explain?

OpenStudy (tkhunny):

Where is the dashed line? Putting it into Intercept Form, \(\dfrac{x}{5}+\dfrac{y}{-10} = 1\), there had better be a broken streak across Quadrant IV with everything away from the Origin shaded.